OpenStudy (anonymous):

need to understand: [ ( ½)^3 – ( ¼ )^2 ] 2^3

OpenStudy (unklerhaukus):

[( ½)^3 – ( ¼ )^2 ] 2^3 =\[ [½ \times ½ \times ½-¼ \times ¼] \times 2 \times 2 \times 2\]

OpenStudy (unklerhaukus):

\[=[1/8-1/16] \times 8\]

OpenStudy (unklerhaukus):

\[=8/16\] =1/2

OpenStudy (anonymous):

the multiplication makes that fraction smaller, got it

OpenStudy (anonymous):

\[((1/2){}^{\wedge}3 - (1/4){}^{\wedge}2)*2{}^{\wedge}3=\frac{1}{2} \]

OpenStudy (anonymous):

that was awesome... i may be back for some real punishment in a bit. fingers crossed